an account offers 4.5% interest compoundd annually. if you deposited $6000 in the account, what would be the account balance at the end of 15 years?
Answer by rfer(10428)

A=P(1+r)^15
A=6000(1.045)^15
A=6000(1.935282443)
A=$11611.69